# An Image is Worth 16x16 Words: Transformers for Image Recognition at Scale

### v1

@@ 1 Introduction
Self-attention-based architectures, in particular Transformers (Vaswani et al., 2017), have become the model of choice in natural language processing (NLP). The dominant approach is to pre-train on a large text corpus and then fine-tune on a smaller task-specific dataset (Devlin et al., 2019). Thanks to Transformers’ computational efficiency and scalability, it has become possible to train models of unprecedented size, with over 100B parameters. With the models and datasets growing, there is still no sign of saturating performance.
@@
When trained on mid-sized datasets such as ImageNet, such models yield modest accuracies of a few percentage points below ResNets of comparable size. This seemingly discouraging outcome may be expected: Transformers lack some of the inductive biases inherent to CNNs, such as translation equivariance and locality, and therefore do not generalize well when trained on insufficient amounts of data.
@@ 2 Related Work
Naive application of self-attention to images would require that each pixel attends to every other pixel. With quadratic cost in the number of pixels, this does not scale to realistic input sizes. Thus, to apply Transformers in the context of image processing, several approximations have been tried in the past: Parmar et al. (2018) applied the self-attention only in local neighborhoods for each query pixel instead of globally. Such local multi-head dot-product self attention blocks can completely replace convolutions (Ramachandran et al., 2019; Cordonnier et al., 2020; Zhao et al., 2020). Alternatively, works such as Sparse Transformers (Child et al., 2019) employ scalable approximations to global self-attention in order to be applicable to images. An alternative way to scale attention is to apply it in blocks of varying sizes (Weissenborn et al., 2019), in the extreme case only along individual axes (Ho et al., 2019; Wang et al., 2020a). Many of these specialized attention architectures demonstrate promising results on computer vision tasks, but require complex engineering to be implemented efficiently on hardware accelerators.
@@
We are not aware of prior application of Transformers with global self-attention to full-sized images. Closest to our model is iGPT (Chen et al., 2020a), which applies Transformers to image pixels after reducing image resolution and color space. The model is trained in an unsupervised fashion as a generative model, and the resulting representation can then be fine-tuned or probed linearly for classification performance, achieving a maximal accuracy of 72% on ImageNet.
@@ 4.1 Setup
For the baseline CNNs, we use ResNet (He et al., 2016), but replace the Batch Normalization layers (Ioffe & Szegedy, 2015) with Group Normalization (Wu & He, 2018), and used standardized convolutions  (Salimans & Kingma, 2016). These modifications improve transfer (Kolesnikov et al., 2020), and we denote the modified model “ResNet (BiT)”. For the hybrids, we feed the intermediate feature maps into ViT with patch size of one “pixel”. To experiment with different sequence lengths, we either (i) take the output of stage 4 of a regular ResNet50 or (ii) remove stage 4, place the same number of layers in stage 3 (keeping the total number of layers), and take the output of this extended stage 3. Option (ii) results in a 4x longer sequence length, and a more expensive ViT model.
@@
Training & Fine-tuning. We train all models, including ResNets, using Adam with $\beta_{1}=0.9$, $\beta_{2}=0.999$, a batch size of 4096 and apply a high weight decay of $0.1$, which we found to be useful for transfer of all models (Appendix D.1 shows that, in contrast to common practices, Adam works slightly better than SGD for ResNets in our setting). We use a linear learning rate warmup and decay, see Appendix B.1 for details. For fine-tuning we use SGD with momentum, batch size 512, for all models, see Appendix B.1.1. For ImageNet results in Table 2, we fine-tuned at higher resolution: $512$ for ViT-L/16 and $518$ for ViT-H/14, and also used Polyak & Juditsky (1992) averaging with a factor of $0.9999$ (Ramachandran et al., 2019; Wang et al., 2020b).
Metrics. We report results on downstream datasets either through few-shot or fine-tuning accuracy. Fine-tuning accuracies capture the performance of each model after fine-tuning it on the respective dataset. Few-shot accuracies are obtained by solving a regularized linear regression problem that maps the (frozen) representation of a subset of training images to $\{-1,1\}^{K}$ target vectors. Though we mainly focus on fine-tuning performance, we sometimes use linear few-shot accuracies for fast on-the-fly evaluation where fine-tuning would be too costly.
@@ 4.2 Comparison to State of the Art
Table 2 shows the results. The smaller ViT-L/16 model pre-trained on JFT-300M outperforms BiT-L (which is pre-trained on the same dataset) on all tasks, while requiring substantially less computational resources to train. The larger model, ViT-H/14, further improves the performance, especially on the more challenging datasets – ImageNet, CIFAR-100, and the VTAB suite. Interestingly, this model still took substantially less compute to pre-train than prior state of the art. However, we note that pre-training efficiency may be affected not only by the architecture choice, but also other parameters, such as training schedule, optimizer, weight decay, etc. We provide a controlled study of performance vs. compute for different architectures in Section 4.4. Finally, the ViT-L/16 model pre-trained on the public ImageNet-21k dataset performs well on most datasets too, while taking fewer resources to pre-train: it could be trained using a standard cloud TPUv3 with 8 cores in approximately 30 days.
@@ 4.3 Pre-training Data Requirements
First, we pre-train ViT models on datasets of increasing size: ImageNet, ImageNet-21k, and JFT-300M. To get the best possible performance on the smaller datasets, we optimize three regularization parameters – weight decay, dropout, and label smoothing. Figure 4 shows the results after fine-tuning to ImageNet (results on other datasets are shown in Table 5) [^2] Note that the ImageNet pre-trained models are also fine-tuned, but again on ImageNet. This is because the resolution increase during fine-tuning improves the performance.. Figure 4 contains the results of fine-tuning on ImageNet. When pre-trained on the smallest dataset, ImageNet, ViT-Large models underperform compared to ViT-Base models, despite heavy regularization. However, with ImageNet-21k pre-training, their performances are similar. Only with JFT-300M, do we see the full benefit of larger models. Figure 4 also shows the performance region spanned by BiT models of different sizes. The BiT CNNs outperform ViT on ImageNet (despite regularization optimization), but with the larger datasets, ViT overtakes.
Second, we train our models on random subsets of 9M, 30M, and 90M as well as the full JFT-300M dataset. We do not perform additional regularization on the smaller subsets and use the same hyper-parameters for all settings. This way, we assess the intrinsic model properties, and not the effect of regularization. We do, however, use early-stopping, and report the best validation accuracy achieved during training. To save compute, we report few-shot linear accuracy instead of full fine-tuning accuracy. Figure 4 contains the results. Vision Transformers overfit more than ResNets with comparable computational cost on smaller datasets. For example, ViT-B/32 is slightly faster than ResNet50; it performs much worse on the 9M subset, but better on 90M+ subsets. The same is true for ResNet152x2 and ViT-L/16. This result reinforces the intuition that the convolutional inductive bias is useful for smaller datasets, but for larger ones, learning the relevant patterns is sufficient, even beneficial.
@@ 4.4 Scaling Study
Figure 5: Performance versus cost for different architectures: Vision Transformers, ResNets, and hybrids. Vision Transformers generally outperform ResNets with the same computational budget. Hybrids improve upon pure Transformers for smaller model sizes, but the gap vanishes for larger models.
@@ 5 Conclusion
We have explored the direct application of Transformers to image recognition. Unlike prior works using self-attention in computer vision, we do not introduce any image-specific inductive biases into the architecture. Instead, we interpret an image as a sequence of patches and process it by a standard Transformer encoder as used in NLP. This simple, yet scalable, strategy works surprisingly well when coupled with pre-training on large datasets. Thus, Vision Transformer matches or exceeds the state of the art on many image classification datasets, whilst being relatively cheap to pre-train.
@@ Acknowledgements
The work was performed in Berlin, Zürich, and Amsterdam. We thank many colleagues at Google for their help, in particular Andreas Steiner for crucial help with the infrastructure and the open-source release of the code; Joan Puigcerver and Maxim Neumann for help with the large-scale training infrastructure; Dmitry Lepikhin, Aravindh Mahendran, Daniel Keysers, Mario Lučić, Noam Shazeer, and Colin Raffel for useful discussions.
@@ Appendix C Additional Results
ImageNet-21k CIFAR-10 98.95 98.79 99.16 99.13 -
CIFAR-100 91.67 91.97 93.44 93.04 -
ImageNet 83.97 81.28 85.15 80.99 -
ImageNet ReaL 88.35 86.63 88.40 85.65 -
Oxford Flowers-102 99.38 99.11 99.61 99.19 -
Oxford-IIIT-Pets 94.43 93.02 94.73 93.09 -
@@
Model Epochs ImageNet ImageNet ReaL CIFAR-10 CIFAR-100 Pets Flowers exaFLOPs
ViT-B/32 7 80.73 86.27 98.61 90.49 93.40 99.27 164
ViT-B/16 7 84.15 88.85 99.00 91.87 95.80 99.56 743
ViT-L/32 7 84.37 88.28 99.19 92.52 95.83 99.45 574
ViT-L/16 7 86.30 89.43 99.38 93.46 96.81 99.66 2586
ViT-L/16 14 87.12 89.99 99.38 94.04 97.11 99.56 5172
ViT-H/14 14 88.08 90.36 99.50 94.71 97.11 99.71 12826
ResNet50x1 7 77.54 84.56 97.67 86.07 91.11 94.26 150
ResNet50x2 7 82.12 87.94 98.29 89.20 93.43 97.02 592
ResNet101x1 7 80.67 87.07 98.48 89.17 94.08 95.95 285
ResNet152x1 7 81.88 87.96 98.82 90.22 94.17 96.94 427
ResNet152x2 7 84.97 89.69 99.06 92.05 95.37 98.62 1681
ResNet152x2 14 85.56 89.89 99.24 91.92 95.75 98.75 3362
ResNet200x3 14 87.22 90.15 99.34 93.53 96.32 99.04 10212
R50x1+ViT-B/32 7 84.90 89.15 99.01 92.24 95.75 99.46 315
R50x1+ViT-B/16 7 85.58 89.65 99.14 92.63 96.65 99.40 855
R50x1+ViT-L/32 7 85.68 89.04 99.24 92.93 96.97 99.43 725
R50x1+ViT-L/16 7 86.60 89.72 99.18 93.64 97.03 99.40 2704
R50x1+ViT-L/16 14 87.12 89.76 99.31 93.89 97.36 99.11 5165
Table 6: Detailed results of model scaling experiments. These correspond to Figure 5 in the main paper.
@@ Appendix D Additional Analyses
ResNets are typically trained with SGD and our use of Adam as optimizer is quite unconventional. Here we show the experiments that motivated this choice. Namely, we compare the fine-tuning performance of two ResNets – 50x1 and 152x2 – pre-trained on JFT with SGD and Adam. For SGD, we use the hyperparameters recommended by Kolesnikov et al. (2020). Results are presented in Table 7. Adam pre-training outperforms SGD pre-training on most datasets and on average. This justifies the choice of Adam as the optimizer used to pre-train ResNets on JFT. Note that the absolute numbers are lower than those reported by Kolesnikov et al. (2020), since we pre-train only for $7$ epochs, not $30$.
@@ D.2 Transformer shape
We ran ablations on scaling different dimensions of the Transformer architecture to find out which are best suited for scaling to very large models. Figure 8 shows 5-shot performance on ImageNet for different configurations. All configurations are based on a ViT model with $8$ layers, $D=1024$, $D_{MLP}=2048$ and a patch size of $32$, the intersection of all lines. We can see that scaling the depth results in the biggest improvements which are clearly visible up until 64 layers. However, diminishing returns are already visible after 16 layers. Interestingly, scaling the width of the network seems to result in the smallest changes. Decreasing the patch size and thus increasing the effective sequence length shows surprisingly robust improvements without introducing parameters. These findings suggest that compute might be a better predictor of performance than the number of parameters, and that scaling should emphasize depth over width if any. Overall, we find that scaling all dimensions proportionally results in robust improvements.
D.3 Positional Embedding

### v2

@@ 1 Introduction
Self-attention-based architectures, in particular Transformers (Vaswani et al., 2017), have become the model of choice in natural language processing (NLP). The dominant approach is to pre-train on a large text corpus and then fine-tune on a smaller task-specific dataset (Devlin et al., 2019). Thanks to Transformers’ computational efficiency and scalability, it has become possible to train models of unprecedented size, with over 100B parameters (Brown et al., 2020; Lepikhin et al., 2020). With the models and datasets growing, there is still no sign of saturating performance.
@@
When trained on mid-sized datasets such as ImageNet without strong regularization, these models yield modest accuracies of a few percentage points below ResNets of comparable size. This seemingly discouraging outcome may be expected: Transformers lack some of the inductive biases inherent to CNNs, such as translation equivariance and locality, and therefore do not generalize well when trained on insufficient amounts of data.
@@ 2 Related Work
Naive application of self-attention to images would require that each pixel attends to every other pixel. With quadratic cost in the number of pixels, this does not scale to realistic input sizes. Thus, to apply Transformers in the context of image processing, several approximations have been tried in the past. Parmar et al. (2018) applied the self-attention only in local neighborhoods for each query pixel instead of globally. Such local multi-head dot-product self attention blocks can completely replace convolutions (Hu et al., 2019; Ramachandran et al., 2019; Zhao et al., 2020). In a different line of work, Sparse Transformers (Child et al., 2019) employ scalable approximations to global self-attention in order to be applicable to images. An alternative way to scale attention is to apply it in blocks of varying sizes (Weissenborn et al., 2019), in the extreme case only along individual axes (Ho et al., 2019; Wang et al., 2020a). Many of these specialized attention architectures demonstrate promising results on computer vision tasks, but require complex engineering to be implemented efficiently on hardware accelerators.
Most related to ours is the model of Cordonnier et al. (2020), which extracts patches of size $2\times 2$ from the input image and applies full self-attention on top. This model is very similar to ViT, but our work goes further to demonstrate that large scale pre-training makes vanilla transformers competitive with (or even better than) state-of-the-art CNNs. Moreover, Cordonnier et al. (2020) use a small patch size of $2\times 2$ pixels, which makes the model applicable only to small-resolution images, while we handle medium-resolution images as well.
@@
Another recent related model is image GPT (iGPT) (Chen et al., 2020a), which applies Transformers to image pixels after reducing image resolution and color space. The model is trained in an unsupervised fashion as a generative model, and the resulting representation can then be fine-tuned or probed linearly for classification performance, achieving a maximal accuracy of 72% on ImageNet.
@@ 3.1 Vision Transformer (ViT)
Inductive bias.
We note that Vision Transformer has much less image-specific inductive bias than CNNs. In CNNs, locality, two-dimensional neighborhood structure, and translation equivariance are baked into each layer throughout the whole model. In ViT, only MLP layers are local and translationally equivariant, while the self-attention layers are global. The two-dimensional neighborhood structure is used very sparingly: in the beginning of the model by cutting the image into patches and at fine-tuning time for adjusting the position embeddings for images of different resolution (as described below). Other than that, the position embeddings at initialization time carry no information about the 2D positions of the patches and all spatial relations between the patches have to be learned from scratch.
@@ 4.1 Setup
For the baseline CNNs, we use ResNet (He et al., 2016), but replace the Batch Normalization layers (Ioffe & Szegedy, 2015) with Group Normalization (Wu & He, 2018), and used standardized convolutions (Qiao et al., 2019). These modifications improve transfer (Kolesnikov et al., 2020), and we denote the modified model “ResNet (BiT)”. For the hybrids, we feed the intermediate feature maps into ViT with patch size of one “pixel”. To experiment with different sequence lengths, we either (i) take the output of stage 4 of a regular ResNet50 or (ii) remove stage 4, place the same number of layers in stage 3 (keeping the total number of layers), and take the output of this extended stage 3. Option (ii) results in a 4x longer sequence length, and a more expensive ViT model.
@@
Training & Fine-tuning. We train all models, including ResNets, using Adam (Kingma & Ba, 2015) with $\beta_{1}=0.9$, $\beta_{2}=0.999$, a batch size of 4096 and apply a high weight decay of $0.1$, which we found to be useful for transfer of all models (Appendix D.1 shows that, in contrast to common practices, Adam works slightly better than SGD for ResNets in our setting). We use a linear learning rate warmup and decay, see Appendix B.1 for details. For fine-tuning we use SGD with momentum, batch size 512, for all models, see Appendix B.1.1. For ImageNet results in Table 2, we fine-tuned at higher resolution: $512$ for ViT-L/16 and $518$ for ViT-H/14, and also used Polyak & Juditsky (1992) averaging with a factor of $0.9999$ (Ramachandran et al., 2019; Wang et al., 2020b).
Metrics. We report results on downstream datasets either through few-shot or fine-tuning accuracy. Fine-tuning accuracies capture the performance of each model after fine-tuning it on the respective dataset. Few-shot accuracies are obtained by solving a regularized least-squares regression problem that maps the (frozen) representation of a subset of training images to $\{-1,1\}^{K}$ target vectors. This formulation allows us to recover the exact solution in closed form. Though we mainly focus on fine-tuning performance, we sometimes use linear few-shot accuracies for fast on-the-fly evaluation where fine-tuning would be too costly.
@@ 4.2 Comparison to State of the Art
Table 2 shows the results. The smaller ViT-L/16 model pre-trained on JFT-300M outperforms BiT-L (which is pre-trained on the same dataset) on all tasks, while requiring substantially less computational resources to train. The larger model, ViT-H/14, further improves the performance, especially on the more challenging datasets – ImageNet, CIFAR-100, and the VTAB suite. Interestingly, this model still took substantially less compute to pre-train than prior state of the art. However, we note that pre-training efficiency may be affected not only by the architecture choice, but also other parameters, such as training schedule, optimizer, weight decay, etc. We provide a controlled study of performance vs. compute for different architectures in Section $4.$ Finally, the ViT-L/16 model pre-trained on the public ImageNet-21k dataset performs well on most datasets too, while taking fewer resources to pre-train: it could be trained using a standard cloud TPUv3 with 8 cores in approximately 30 days.
@@ 4.3 Pre-training Data Requirements
First, we pre-train ViT models on datasets of increasing size: ImageNet, ImageNet-21k, and JFT-300M. To boost the performance on the smaller datasets, we optimize three basic regularization parameters – weight decay, dropout, and label smoothing. Figure 4 shows the results after fine-tuning to ImageNet (results on other datasets are shown in Table 5) [^2] Note that the ImageNet pre-trained models are also fine-tuned, but again on ImageNet. This is because the resolution increase during fine-tuning improves the performance.. When pre-trained on the smallest dataset, ImageNet, ViT-Large models underperform compared to ViT-Base models, despite (moderate) regularization. With ImageNet-21k pre-training, their performances are similar. Only with JFT-300M, do we see the full benefit of larger models. Figure 4 also shows the performance region spanned by BiT models of different sizes. The BiT CNNs outperform ViT on ImageNet, but with the larger datasets, ViT overtakes.
Second, we train our models on random subsets of 9M, 30M, and 90M as well as the full JFT-300M dataset. We do not perform additional regularization on the smaller subsets and use the same hyper-parameters for all settings. This way, we assess the intrinsic model properties, and not the effect of regularization. We do, however, use early-stopping, and report the best validation accuracy achieved during training. To save compute, we report few-shot linear accuracy instead of full fine-tuning accuracy. Figure 4 contains the results. Vision Transformers overfit more than ResNets with comparable computational cost on smaller datasets. For example, ViT-B/32 is slightly faster than ResNet50; it performs much worse on the 9M subset, but better on 90M+ subsets. The same is true for ResNet152x2 and ViT-L/16. This result reinforces the intuition that the convolutional inductive bias is useful for smaller datasets, but for larger ones, learning the relevant patterns directly from data is sufficient, even beneficial.
@@
Figure 5: Performance versus pre-training compute for different architectures: Vision Transformers, ResNets, and hybrids. Vision Transformers generally outperform ResNets with the same computational budget. Hybrids improve upon pure Transformers for smaller model sizes, but the gap vanishes for larger models.
@@ 5 Conclusion
We have explored the direct application of Transformers to image recognition. Unlike prior works using self-attention in computer vision, we do not introduce image-specific inductive biases into the architecture apart from the initial patch extraction step. Instead, we interpret an image as a sequence of patches and process it by a standard Transformer encoder as used in NLP. This simple, yet scalable, strategy works surprisingly well when coupled with pre-training on large datasets. Thus, Vision Transformer matches or exceeds the state of the art on many image classification datasets, whilst being relatively cheap to pre-train.
@@ Acknowledgements
The work was performed in Berlin, Zürich, and Amsterdam. We thank many colleagues at Google for their help, in particular Andreas Steiner for crucial help with the infrastructure and the open-source release of the code; Joan Puigcerver and Maxim Neumann for help with the large-scale training infrastructure; Dmitry Lepikhin, Aravindh Mahendran, Daniel Keysers, Mario Lučić, Noam Shazeer, Ashish Vaswani, and Colin Raffel for useful discussions.
@@ Appendix C Additional Results
ImageNet-21k CIFAR-10 98.95 98.79 99.16 99.13 99.27
CIFAR-100 91.67 91.97 93.44 93.04 93.82
ImageNet 83.97 81.28 85.15 80.99 85.13
ImageNet ReaL 88.35 86.63 88.40 85.65 88.70
Oxford Flowers-102 99.38 99.11 99.61 99.19 99.51
Oxford-IIIT-Pets 94.43 93.02 94.73 93.09 94.82
@@
Epochs ImageNet ImageNet ReaL CIFAR-10 CIFAR-100 Pets Flowers exaFLOPs
name
ViT-B/32 7 80.73 86.27 98.61 90.49 93.40 99.27 55
ViT-B/16 7 84.15 88.85 99.00 91.87 95.80 99.56 224
ViT-L/32 7 84.37 88.28 99.19 92.52 95.83 99.45 196
ViT-L/16 7 86.30 89.43 99.38 93.46 96.81 99.66 783
ViT-L/16 14 87.12 89.99 99.38 94.04 97.11 99.56 1567
ViT-H/14 14 88.08 90.36 99.50 94.71 97.11 99.71 4262
ResNet50x1 7 77.54 84.56 97.67 86.07 91.11 94.26 50
ResNet50x2 7 82.12 87.94 98.29 89.20 93.43 97.02 199
ResNet101x1 7 80.67 87.07 98.48 89.17 94.08 95.95 96
ResNet152x1 7 81.88 87.96 98.82 90.22 94.17 96.94 141
ResNet152x2 7 84.97 89.69 99.06 92.05 95.37 98.62 563
ResNet152x2 14 85.56 89.89 99.24 91.92 95.75 98.75 1126
ResNet200x3 14 87.22 90.15 99.34 93.53 96.32 99.04 3306
R50x1+ViT-B/32 7 84.90 89.15 99.01 92.24 95.75 99.46 106
R50x1+ViT-B/16 7 85.58 89.65 99.14 92.63 96.65 99.40 274
R50x1+ViT-L/32 7 85.68 89.04 99.24 92.93 96.97 99.43 246
R50x1+ViT-L/16 7 86.60 89.72 99.18 93.64 97.03 99.40 859
R50x1+ViT-L/16 14 87.12 89.76 99.31 93.89 97.36 99.11 1668
Table 6: Detailed results of model scaling experiments. These correspond to Figure 5 in the main paper. We show transfer accuracy on several datasets, as well as the pre-training compute (in exaFLOPs).
@@ Appendix D Additional Analyses
ResNets are typically trained with SGD and our use of Adam as optimizer is quite unconventional. Here we show the experiments that motivated this choice. Namely, we compare the fine-tuning performance of two ResNets – 50x1 and 152x2 – pre-trained on JFT with SGD and Adam. For SGD, we use the hyperparameters recommended by Kolesnikov et al. (2020). Results are presented in Table 7. Adam pre-training outperforms SGD pre-training on most datasets and on average. This justifies the choice of Adam as the optimizer used to pre-train ResNets on JFT. Note that the absolute numbers are lower than those reported by Kolesnikov et al. (2020), since we pre-train only for 77 epochs, not $30$.
@@ D.2 Transformer shape
We ran ablations on scaling different dimensions of the Transformer architecture to find out which are best suited for scaling to very large models. Figure 8 shows 5-shot performance on ImageNet for different configurations. All configurations are based on a ViT model with 88 layers, $D=1024$, $D_{MLP}=2048$ and a patch size of $32$, the intersection of all lines. We can see that scaling the depth results in the biggest improvements which are clearly visible up until 64 layers. However, diminishing returns are already visible after 16 layers. Interestingly, scaling the width of the network seems to result in the smallest changes. Decreasing the patch size and thus increasing the effective sequence length shows surprisingly robust improvements without introducing parameters. These findings suggest that compute might be a better predictor of performance than the number of parameters, and that scaling should emphasize depth over width if any. Overall, we find that scaling all dimensions proportionally results in robust improvements.
Figure 9: Comparison of class-token and global average pooling classifiers. Both work similarly well, but require different learning-rates.
D.3 Head Type and class token
In order to stay as close as possible to the original Transformer model, we made use of an additional [class] token, which is taken as image representation. The output of this token is then transformed into a class prediction via a small multi-layer perceptron (MLP) with $\tanh$ as non-linearity in the single hidden layer.
This design is inherited from the Transformer model for text, and we use it throughout the main paper. An initial attempt at using only image-patch embeddings, globally average-pooling (GAP) them, followed by a linear classifier—just like ResNet’s final feature map—performed very poorly. However, we found that this is neither due to the extra token, nor to the GAP operation. Instead, the difference in performance is fully explained by the requirement for a different learning-rate, see Figure 9.
D.4 Positional Embedding